OpenStudy (anonymous):

Jacob spends 60 minutes in the gym every day doing freehand exercises and running on the treadmill. He spends 30 minutes more running on the treadmill than doing freehand exercises. Part A: Write a pair of linear equations to show the relationship between the number of minutes Jacob does freehand exercises (y) and the number of minutes he runs on the treadmill (x). (5 points) Part B: How much time does Jacob spend on doing freehand exercises? (3 points) Part C: Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. (2 points)

OpenStudy (anonymous):

@confluxepic @ganeshie8 @HexiFlexiDexi @helpkay @HarleyQuinn98 @YanaSidlinskiy @yrdnfl @RaphaelFilgueiras @rsadhvika @Diyadiya @flower2468

OpenStudy (flower2468):

what do you think the answer is

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

this is not multiple choice by the way :)

OpenStudy (flower2468):

I know, but do you have any work done or do just don't know where to start?

OpenStudy (anonymous):

i dont know where to start

OpenStudy (anonymous):

how do i do this?

OpenStudy (flower2468):

well a linear equation is:A linear equation is an algebraic equation in which each term is either a constant or the product of a constant and (the first power of) a single variable.

OpenStudy (flower2468):

@Qwertty123

OpenStudy (qwertty123):

Hey! :D Part A x + y = 60 x = 30 + y

OpenStudy (qwertty123):

So for B...what do you get?

OpenStudy (anonymous):

let me try this hold on this may take a second

OpenStudy (anonymous):

do i solve this by substitution?

OpenStudy (jameshorton):

y = freehand and x = treadmill y + x = 60 x = y + 30 y + y + 30 = 60 2y = 60 - 30 2y = 30 y = 15 freehand is 15 minutes x + y = 60 x + 15 = 60 x = 60 - 15 x = 45 yes....he can spend up to 45 minutes doing treadmill

OpenStudy (qwertty123):

Wait I thought B was 15 minutes ? Did I do it wrong?

OpenStudy (texaschic101):

b is 15 minutes

OpenStudy (qwertty123):

oh okay :D

OpenStudy (anonymous):

so how do i do part c?

OpenStudy (qwertty123):

For Part C.... From what you found above... What do you think.. Yes or no?

OpenStudy (anonymous):

I think yes but it says to explain your reason. so do i say something like "yes he could of done 40 minutes on the teadmill because the equation said that he did 45" ?

OpenStudy (texaschic101):

yes

OpenStudy (qwertty123):

C. Yes, the time spent in the gym would be 55 minutes(15+40).

OpenStudy (anonymous):

thank you!!

OpenStudy (qwertty123):

You Welcome! :) See you around.